Find the slope-intercept form for the line satisfying the following conditions.
x-intercept 2, y-intercept 1/3
The slope-intercept form for the line is?

Answer by lwsshak3(6770)

Find the slope-intercept form for the line satisfying the following conditions.
x-intercept 2, y-intercept 1/3
The slope-intercept form for the line is?
**
note: two points are given: (2,0) and (0,1/3)
slope: ∆y/∆x=(1/3-0)/(0-2)=(1/3)/-2=-1/6
equation: y=-x/6+b
solve for b using coordinates of one of given points (2,0)
0=2*(-1/6)+b
0=-2/6+b
b=2/6=1/3
slope-intercept form for the line: y=-x/6+1/3